# Tracial states on groupoid $C^{*}$-algebras and essential freeness

### v2

@@ 5.2. Tracial ideals and quasidiagonal traces
According to [5, Proposition 5.4], it suffices to show that
@@
A question of N. Brown asks whether every amenable tracial state is quasidiagonal (see [6, Question 6.7(2)]). Substantial progress on this question has recently been made in [13, 32] as follows:
@@
Let $\mathcal{G}$ be a locally compact, Hausdorff, second countable and étale groupoid such that $C_{r}^{*}(\mathcal{G})$ is an exact $C^{*}$-algebra and $\mathcal{G}$ satisfies the strong Baum-Connes conjecture with all coefficients in the sense of [9, Definition 3.6]. If $\mathcal{G}$ is also essentially free, then every amenable tracial state on $C_{r}^{*}(\mathcal{G})$ is quasidiagonal.
@@
Let $\tau$ be any amenable tracial state on $C_{r}^{*}(\mathcal{G})$. Then $\tau$ vanishes on the tracial ideal $I_{\tau}$ by Remark 5.10, and hence there is an induced faithful tracial state $\dot{\tau}$ on $C_{r}^{*}(\mathcal{G})/I_{\tau}$ such that $\tau=\dot{\tau}\circ p$, where $p:C_{r}^{*}(\mathcal{G})\rightarrow C_{r}^{*}(\mathcal{G})/I_{\tau}$ is the quotient map. Clearly, the quasidiagonality of $\tau$ follows from the fact that $\dot{\tau}$ is quasidiagonal.
Since $\tau$ is amenable and $C_{r}^{*}(\mathcal{G})$ is exact, it follows that $\dot{\tau}$ is also amenable (see [7, Proposition 6.3.5 (4)]). By Corollary 5.12, there is a closed invariant subset $D$ of $\mathcal{G}^{(0)}$ such that $C_{r}^{*}(\mathcal{G})/I_{\tau}\cong C_{r}^{*}(\mathcal{G}_{D})$, which is exact as quotients of exact $C^{*}$-algebras are exact by [7, Theorem 10.2.5]. As $\mathcal{G}$ is second countable and satisfies the strong Baum-Connes conjecture with all coefficients, $C_{r}^{*}(\mathcal{G}_{D})$ is separable and satisfies the UCT by [9, Theorem 4.11 and Corollary 4.2]. Thus, we conclude the quasidiagonality of $\dot{\tau}$ from Theorem 5.13. ∎
@@
Let $\mathcal{G}$ be a locally compact, Hausdorff, second countable, amenable and étale groupoid, which is also essentially free. Then every tracial state on $C_{r}^{*}(\mathcal{G})$ is quasidiagonal.
@@
If $G$ is amenable, then $\mathcal{G}$ satisfies the strong Baum-Connes conjecture with all coefficients by [9, Corollary 3.15] and $C_{r}^{*}(G)$ is nuclear (hence also exact). As every tracial state on a nuclear $C^{*}$-algebra is amenable [7, Proposition 6.3.4], we complete the proof by Theorem 5.14. ∎
Acknowledgement: We wish to thank Johannes Christensen, Sergey Neshveyev and Eduardo Scarparo for helpful discussions on this topic.

### v3

@@ 5.2. Tracial ideals and quasidiagonal traces
Firstly, we note that $\mu$ being invariant implies that $\mathrm{supp}\mu$ is invariant. According to [5, Proposition 5.4], it suffices to show that
@@
A question of N. Brown asks whether every amenable tracial state is quasidiagonal (see [6, Question 6.7(2)]) in the following sense:
Definition 5.13.
A tracial state $\tau$ on a $C^{*}$-algebra $A$ is quasidiagonal if there is a net of contractive completely positive maps $\phi_{i}:A\rightarrow M_{k(i)}$ such that
  • $\tau(a)=\lim_{i}\text{tr}\circ\phi_{i}(a)$ for all $a\in A$;
  • $\lim_{i}||\phi_{i}(ab)-\phi_{i}(a)\phi_{i}(b)||=0$ for all $a,b\in A$.
Substantial progress on this question has recently been made in [13, 33] as follows:
@@
Let $\mathcal{G}$ be a locally compact, Hausdorff, second-countable and étale groupoid such that $C_{r}^{*}(\mathcal{G})$ is an exact $C^{*}$-algebra and $\mathcal{G}$ satisfies the strong Baum–Connes conjecture with all coefficients in the sense of [9, Definition 3.6]. If $\mathcal{G}$ is also essentially free, then every amenable tracial state on $C_{r}^{*}(\mathcal{G})$ is quasidiagonal.
@@
Let $\tau$ be any amenable tracial state on $C_{r}^{*}(\mathcal{G})$. Then $\tau$ vanishes on the tracial ideal $I_{\tau}$ by Remark 5.10, and hence there is an induced faithful tracial state $\dot{\tau}$ on $C_{r}^{*}(\mathcal{G})/I_{\tau}$ such that $\tau=\dot{\tau}\circ p$, where $p:C_{r}^{*}(\mathcal{G})\rightarrow C_{r}^{*}(\mathcal{G})/I_{\tau}$ is the quotient map. To see that $\tau$ is quasidiagonal, it suffices to show that $\dot{\tau}$ is quasidiagonal.
Since $\tau$ is amenable and $C_{r}^{*}(\mathcal{G})$ is exact, it follows that $\dot{\tau}$ is also amenable (see [7, Proposition 6.3.5 (4)]). By Corollary 5.12, there is a closed invariant subset $D$ of $\mathcal{G}^{(0)}$ such that $C_{r}^{*}(\mathcal{G})/I_{\tau}\cong C_{r}^{*}(\mathcal{G}_{D})$, which is exact as quotients of exact $C^{*}$-algebras are exact by [7, Theorem 10.2.5]. As $\mathcal{G}$ is second-countable and satisfies the strong Baum–Connes conjecture with all coefficients, $C_{r}^{*}(\mathcal{G}_{D})$ is separable and satisfies the UCT by [9, Theorem 4.11 and Corollary 4.2]. Thus, we conclude the quasidiagonality of $\dot{\tau}$ from Theorem 5.14. ∎
@@
Let $\mathcal{G}$ be a locally compact, Hausdorff, second-countable, amenable and étale groupoid, which is also essentially free. Then every tracial state on $C_{r}^{*}(\mathcal{G})$ is quasidiagonal.
@@
If $G$ is amenable, then $\mathcal{G}$ satisfies the strong Baum–Connes conjecture with all coefficients by [9, Corollary 3.15] and $C_{r}^{*}(G)$ is nuclear (hence also exact). As every tracial state on a nuclear $C^{*}$-algebra is amenable [7, Proposition 6.3.4], we complete the proof by Theorem 5.15. ∎
Acknowledgement: We wish to thank Johannes Christensen, Sergey Neshveyev and Eduardo Scarparo for helpful discussions on this topic. Finally, we would like to thank the anonymous reviewers for the helpful comments.